Jay has $320.00. She gives .4 of it to Ray. She gives 025 of it to boy. She keeps the rest. How much does each person get
Found 2 solutions by jorel555, 428225:
Answer by jorel555(1290)

If he has $320 to start, then:
320*.4=$128 to Ray, and
320*.025=$8 to boy!!!!!!!!!!!!!

Answer by 428225(90)

I am guessing that the question should say "She gives 0.25 of it to [a] boy". If so, I shall proceed to answer your question.

To figure out how much Ray got you can simply calculate:

320 * 0.4 = ?

To figure out how much the boy got you can calculate:

320 * 0.25 = ?

To figure out how much Jay got you can calculate:

320 * (1-0.25-0.4) = ?